OpenStudy (anonymous):

Rebekah purchased a sweater that was reduced by twenty-five percent (or one-fourth) of its original price. She also used a store coupon for five dollars. When she made the purchase, she paid the cashier nineteen dollars. What was the original price of the sweater?

OpenStudy (lukebluefive):

I don't think you even gave the correct equation to solve for x. The sweater's price was reduced by 25%, not to 25% of the original price.

jigglypuff314 (jigglypuff314):

ah right, thanks @LukeBlueFive

jigglypuff314 (jigglypuff314):

1 ( the original price) - 0.25 (discount) = 0.75 (new price) so 0.75x - 5 = 19